OpenStudy (anonymous):

Complete the square to solve the equation. x^2+x=7/4

OpenStudy (anonymous):

divide the number multiplying x (here it is 1) by 2: (x + 1/2)^2 expanding this gives a last term of 1/4 so to get back to x^2+x we need to subtract 1/4 so (x + 1)^2 -1/4 = 7/4 (x + 1)^2 = 2 x+1 = sqrt2 or -sqrt1 x = sqrt2 - 1 , -sqrt2 - 1

OpenStudy (anonymous):

these work out to 0.4142 , -2.4142

OpenStudy (anonymous):

Okay thanks again!

OpenStudy (anonymous):

the aim is to get a perfect square on left side and then take square roots of both sides

OpenStudy (anonymous):

yw